The width of a rectangular lawn is 3m less than half its length. The perimeter is 42m. Show that w+l=21. write an equation for the width and calculate the width of the lawn.
Answer by nerdybill(7384)

The width of a rectangular lawn is 3m less than half its length. The perimeter is 42m. Show that w+l=21. write an equation for the width and calculate the width of the lawn.
.
Let W = width
and L = length
.
2(W+L) = 42
dividing both sides by 2 we get answer to first part:
W+L = 21
.
From: "width of a rectangular lawn is 3m less than half its length" we get:
width = L/2 - 3
.
Plug above into:
W+L = 21
L/2 - 3 +L = 21
L - 6 +2L = 42
-6 +3L = 42
3L = 48
L = 16m (length)
.
calculating width:
width = L/2 - 3
width = 16/2 - 3
width = 8 - 3
width = 5m (width)